Physics 152—Accelerated Physics II: Electricity, Magnetism, and Optics Homework Assignment #9 Due Friday, November 8, 2024, 9:30xam

Mastering Physics

Do the on-line assignment “HW #09.”

Pencil-and-Paper Problems There are no pencil and paper problems this week.

Academic Honesty Since there is no written work this week, there is no place for you to write down and acknowledge any collaborators. Nevertheless, the same basic principles of intellectual honesty apply. You may get help from others, but the work you ultimately turn in as your own should be your own.